# Optimized Minimal 3D Gaussian Splatting

### v1

@@ 3.1 Overall Architecture
3.2 Sub-Vector Quantization
Vector quantization (VQ) [18] has shown high efficiency for representing Gaussian attributes, capitalizing on their inherently vectorized structure and strong global coherence across an entire scene. However, to maintain high fidelity, a large codebook size is required, inevitably resulting in substantial computational overhead and increased training complexity [60] (Fig. 3(a)). To address these issues, Residual Vector Quantization (R-VQ) [60] has been used as a hierarchical quantization strategy [29], progressively refining representations while reducing the size of each individual codebook. However, as shown in Fig. 3(b), multiple code indices per attribute result in increased storage overhead, illustrating a tradeoff between reducing per-codebook complexity and increasing overall storage requirements.
To navigate this tradeoff, we propose Sub-Vector Quantization (SVQ), which partitions the attribute vector into multiple sub-vectors and applies vector quantization separately to each component (Fig. 3(c)), motivated by Product Quantization [24]. By reducing the dimensionality of each quantized unit, SVQ allows for smaller codebooks and more efficient lookups, which can balance computational cost and storage efficiency while maintaining high fidelity. We can apply SVQ to an input vector $z\in\mathbb{R}^{ML}$, where $M$ and $L$ represent the total number of sub-vectors (partitions) and the sub-vector length, respectively. Each partition $m\in\{1,...,M\}$ has an independent codebook $C^{(m)}\in\mathbb{R}^{B\times L}$, where $B$ denotes the number of codewords per codebook. The codeword selection is based on the nearest match from $C^{(m)}$, with $C^{(m)}[j]$ representing the $j$-th codeword corresponding to the $m$-th sub-vector. More formally, SVQ-applied vector $\hat{z}$ can be formulated as follows,
$i_{m}=\arg\min_{j}\|z_{m}-C^{(m)}[j]\|^{2}_{2},\quad m\in\{1,...,M\},$ (6)
SVQ ensures significantly reduced computation with small codebooks compared to VQ. We apply SVQ to geometric attributes $s_{n},r_{n}$, resulting in quantized vectors $\hat{s}_{n},\hat{r}_{n}$, which are then used for 3DGS rendering. For the appearance features $T_{n},V_{n}$, we first concatenate them and apply SVQ. The resulting quantized features are then split back into two components $\hat{T}_{n},\hat{V}_{n}$, which replace $T_{n}$ and $V_{n}$ in Eqs. 3 and 4.
Table 1: Quantitative results of OMG evaluated on the Mip-NeRF 360 dataset. Baseline results are sourced from the LocoGS [50] paper, where the rendering results were obtained using an NVIDIA RTX 3090 GPU. Our rendering performance was measured using the same GPU, with the values in parentheses obtained from an NVIDIA RTX 4090 GPU. We highlight the results among compression methods by coloring the best, second-best, and third-best performances.
Mip-NeRF 360
3DGS 27.44 0.813 0.218 822.6 127
Scaffold-GS [34] 27.66 0.812 0.223 187.3 122
CompGS [41] 27.04 0.804 0.243 22.93 236
Compact-3DGS [29] 26.95 0.797 0.244 26.31 143
C3DGS [42] 27.09 0.802 0.237 29.98 134
LightGaussian [11] 26.90 0.800 0.240 53.96 244
EAGLES [17] 27.10 0.807 0.234 59.49 155
SOG [39] 27.01 0.800 0.226 43.77 134
HAC [7] 27.49 0.807 0.236 16.95 110
LocoGS-S [50] 27.04 0.806 0.232 7.90 310
LocoGS-L [50] 27.33 0.814 0.219 13.89 270
OMG-XS 27.06 0.807 0.243 4.06 350 (612)
OMG-M 27.21 0.814 0.229 5.31 298 (511)
OMG-XL 27.34 0.819 0.218 6.82 251 (416)
3.3 Local Distinctiveness for Important Scoring
OMG adopts importance scoring to identify essential Gaussians and retain a minimal number of them. Existing scoring-based pruning methods typically determine the importance of each Gaussian based on its blending weights (the values multiplied by $c_{k}$ in Eq. 1) across training-view renderings. We use two factors as our baseline metric: (1) whether it has been the most dominant contributor for at least one ray [12, 13] and (2) its total blending weight contribution across all training rays [43, 17]. Formally, we define the base importance score $\bar{I}$ as:
To mitigate these issues, we propose incorporating a local distinctiveness metric into the importance computation. Specifically, we introduce an additional term that measures the appearance (static) feature similarity among neighboring Gaussians, ensuring that distinct Gaussians show high importance. The final importance score is defined as:
$I_{i}=\bar{I}_{i}\left(\frac{1}{K}\sum_{j\in\mathcal{N}_{i}^{K}}\|T_{i}-T_{j}\|_{1}\right)^{\lambda},$ (8)
where $\mathcal{N}_{i}^{K}$ denotes the set of $K$-nearest neighbors of Gaussian $i$ and $\lambda$ is a scaling factor that adjusts the sensitivity to appearance variation. As computing exact $K$-nearest neighbors for every Gaussian is computationally expensive, we approximate neighbor selection by sorting Gaussians in Morton order and selecting Gaussians with adjacent indices as their local neighbors. We remove low-importance Gaussians using CDF-based thresholding [30] with a threshold $\tau$.
Table 2: Quantitative results of OMG evaluated on the Tanks&Temples and Deep Blending datasets. Baseline results are sourced from the LocoGS [50] paper, where the rendering results were obtained using an NVIDIA RTX 3090 GPU. Our rendering performance was measured using the same GPU, with the values in parentheses obtained from an NVIDIA RTX 4090 GPU.
@@ 3.3 Local Distinctiveness for Important Scoring
3DGS [26] 23.67 0.844 0.179 452.4 175 29.48 0.900 0.246 692.5 134
Scaffold-GS  [34] 24.11 0.855 0.165 154.3 109 30.28 0.907 0.243 121.2 194
CompGS [41] 23.29 0.835 0.201 14.23 329 29.89 0.907 0.253 15.15 301
Compact-3DGS [29] 23.33 0.831 0.202 18.97 199 29.71 0.901 0.257 21.75 184
C3DGS [42] 23.52 0.837 0.188 18.58 166 29.53 0.899 0.254 24.96 143
LightGaussian [11] 23.32 0.829 0.204 29.94 379 29.12 0.895 0.262 45.25 287
EAGLES [17] 23.14 0.833 0.203 30.18 244 29.72 0.906 0.249 54.45 137
SOG [39] 23.54 0.833 0.188 24.42 222 29.21 0.891 0.271 19.32 224
HAC [7] 24.08 0.846 0.186 8.42 129 29.99 0.902 0.268 4.51 235
LocoGS-S [50] 23.63 0.847 0.169 6.59 333 30.06 0.904 0.249 7.64 334
OMG-M 23.52 0.842 0.189 3.22 555 (887) 29.77 0.908 0.253 4.34 524 (894)
OMG-L 23.60 0.846 0.181 3.93 478 (770) 29.88 0.910 0.247 5.21 479 (810)
4 Experiment
4.1 Implementation Details
Following the previous works, we evaluated our approach on three real-world datasets, Mip-NeRF 360 [2], Tanks&Temples [28], and Deep Blending [21]. Our model is implemented upon Mini-Splatting [12], one of the methods achieving high performance with a small number of Gaussians. We have conducted simple post-processings after training:
  • Applying 16-bit quantization to the position and compressing with G-PCC [48].
  • Huffman encoding [23] to SVQ indices.
  • Storing all the components into a single file with LZMA [1] compression.
We provide five OMG variants (XS, X, M, L, XL), adjusting storage requirements. The only factor controlling the storage is the CDF-based threshold value $\tau$ of Gaussian importance, which is set to 0.96, 0.98, 0.99, 0.999, and 0.9999 for each variant, respectively. Further implementation details are provided in the supplementary materials.
@@ 4.2 Performance Evaluation
Compression performance. Tabs. 1 and 2 compare the performance of OMG against various baseline methods on the Mip-NeRF 360, Tanks & Temples, and DeepBlending datasets. OMG consistently shows the smallest storage requirements while maintaining high performance across all datasets, achieving state-of-the-art (SOTA) results. Notably, on the Mip-NeRF 360 dataset, OMG-XS achieves nearly a 50% reduction in storage compared to the small variant of LocoGS [50], the previous SOTA compression method, while retaining PSNR and SSIM. With over 30% reduced storage, OMG-M outperforms LocoGS-S in all quality metrics. Moreover, OMG-XL surpasses LocoGS-L in all metrics, even though requiring less storage than LocoGS-S.
The qualitative results presented in Fig. 4 also demonstrate the strong performance of OMG. Despite achieving over 100× compression compared to 3DGS, OMG maintains comparable visual quality. Especially, in the bicycle scene, OMG-XS achieves over 300× compression relative to 3DGS while accurately reconstructing details that 3DGS fails to represent, resulting in a blurry area (highlighted in red) in its rendering. This superiority can be attributed to the blur split technique of our baseline model, mini-splatting [12]. Despite reducing the number of Gaussians by an additional 20% compared to mini-splatting, OMG-XS retains high visual fidelity, demonstrating its effectiveness in extreme compression scenarios.
Computational efficiency. OMG achieves remarkable efficiency alongside high performance. As shown in Tab. 3, OMG shows superior scene fidelity with significantly fewer Gaussian primitives compared to LocoGS. This reduction results in substantial rendering speed improvements of 13%, 67%, and 57% for the Mip-NeRF 360, Tank&Temples, and Deep Blending datasets (Tabs. 1 and 2), respectively, compared to LocoGS, highlighting its potential for real-time rendering on low-capacity devices. Furthermore, OMG accelerates training speed. The substantial improvement over LocoGS can be attributed to two key factors: the reduced number of Gaussians and the absence of a large neural field. By efficiently exploiting coarse spatial information through a tiny MLP, OMG achieves high computational efficiency.
Figure 4: Qualitative results of OMG compared to 3DGS. We provide per-image rendering PSNR with storage requirements for each scene.
Table 3: Efficiency comparison of OMG variants compared to LocoGS evaluated on the Mip-NeRF 360 dataset. We present training time, the number of Gaussians, and the storage requirement with rendering quality.
@@
LocoGS-S $\approx$1h 1.09M 7.9 27.04 0.806 0.232
@@
Table 4: Ablation study of OMG using the Mip-NeRF 360 dataset. We evaluate the contribution of the space feature integration and local distinctiveness (LD) scoring.
@@
Table 5: Ablation study on SVQ using the Mip-NeRF 360 dataset. We substitute SVQ to VQ.
Method Training #Gauss Size PSNR SSIM LPIPS
OMG-XS 20m 15s 427K 4.06 27.06 0.807 0.243
SVQ $\rightarrow$ VQ 21m 22s 426K 3.99 26.97 0.805 0.245
@@ 4.3 Ablation Study
OMG architecture. OMG leverages a highly compact neural field to capture coarse spatial information while reducing the number of learnable parameters per Gaussian. Tab. 4 validates the contribution of this space feature. Although the total number of Gaussians slightly increases, performance significantly degrades. The absence of spatial information introduces instability in attribute learning, hindering effective importance scoring. This trend is consistently observed in both our small and medium models, highlighting the effectiveness of the space feature despite its minimal parameter overhead.
@@
Local distinctiveness scoring. OMG improves Gaussian pruning by incorporating local distinctiveness (LD) scoring into the importance estimation. Tab. 3 evaluates the impact of LD scoring, showing that despite a similar number of Gaussians, the inclusion of LD scoring leads to a significant performance improvement. This effect becomes even more pronounced when the target Gaussian number is lower, demonstrating that LD scoring provides an effective approach for further reducing a sparse set of Gaussians. Furthermore, when both the space feature and LD scoring are removed, the model experiences the most substantial performance drop. This indicates that the two contributions are orthogonal, independently contributing to model efficiency and performance.
Sub-vector quantization. To assess the impact of SVQ, we replaced it with VQ while maintaining the proposed clever training strategy, where K-means is performed only once before the final 1K iterations, after which only the codebook is updated. Despite this adjustment, VQ leads to a 5% increase in training time, entirely due to K-means initialization. This is because large codebooks are inevitably required for accurate representation (here, we use $2^{14}$ codes for scale and rotation and $2^{13}$ codes for each appearance feature). Nevertheless, VQ results in a decline in rendering quality, as shown in Tab. 5. To improve rendering quality, VQ requires an enlarged codebook, which is constrained by significant computational and memory overhead. Moreover, the increase in codebook size reduces storage efficiency when the number of Gaussians is small. In contrast, SVQ offers scalability of representation by flexibly adjusting the bit budget with sub-vector length in regard to the capacity-performance trade-off.

### v2

@@ 4.1 Implementation details
LocoGS* (COLMAP)
27.09 0.798 0.250 7.96 (396)
[…]
Table 2: Quantitative results of OMG evaluated on the Tanks&Temples and Deep Blending datasets. Baseline results are sourced from the LocoGS [locogs] paper, where the rendering results were obtained using an NVIDIA RTX 3090 GPU. Our rendering performance was measured using the same GPU, with the values in parentheses obtained from an NVIDIA RTX 4090 GPU.
@@
3DGS [3dgs] 23.67 0.844 0.179 452.4 175 29.48 0.900 0.246 692.5 134
Scaffold-GS  [scaffoldgs] 24.11 0.855 0.165 154.3 109 30.28 0.907 0.243 121.2 194
Mini-Splatting [msv1] 23.41 0.846 0.180 67.6 (1095) 30.04 0.910 0.241 124.9 (902)
[…]
Compact-3DGS [c3dgs] 23.33 0.831 0.202 18.97 199 29.71 0.901 0.257 21.75 184
C3DGS [morton3d] 23.52 0.837 0.188 18.58 166 29.53 0.899 0.254 24.96 143
[…]
SOG [c3d2d] 23.54 0.833 0.188 24.42 222 29.21 0.891 0.271 19.32 224
[…]
@@ 4.2 Performance evaluation
Compression performance. Tables˜1 and 2 compare the performance of OMG against various baseline methods on the Mip-NeRF 360, Tanks & Temples, and DeepBlending datasets. OMG consistently shows the smallest storage requirements while maintaining high performance across all datasets, achieving state-of-the-art (SOTA) results. Notably, on the Mip-NeRF 360 dataset, OMG-XS achieves nearly a 50% reduction in storage compared to the small variant of LocoGS [locogs], the previous SOTA compression method, while retaining PSNR and SSIM. With over 30% reduced storage, OMG-M outperforms LocoGS-S in all quality metrics. Moreover, OMG-XL surpasses LocoGS-L in all metrics, even though requiring less storage than LocoGS-S.
The qualitative results presented in Figure˜4 also demonstrate the strong performance of OMG. Despite achieving over 100× compression compared to 3DGS, OMG maintains comparable visual quality. Especially, in the bicycle scene, OMG-XS achieves over 300× compression relative to 3DGS while accurately reconstructing details that 3DGS fails to represent, resulting in a blurry area (highlighted in red) in its rendering. This superiority can be attributed to the blur split technique of our baseline model, Mini-Splatting [msv1]. Despite reducing the number of Gaussians by an additional 20% compared to Mini-Splatting (Table˜3), OMG-XS retains high visual fidelity, demonstrating its effectiveness in extreme compression scenarios.
Computational efficiency. OMG achieves remarkable efficiency alongside high performance. As shown in Table˜3, OMG shows superior scene fidelity with significantly fewer Gaussian primitives compared to LocoGS. This reduction results in substantial rendering speed improvements of 13%, 67%, and 57% for the Mip-NeRF 360, Tank&Temples, and Deep Blending datasets (Tables˜1 and 2), respectively, compared to LocoGS, highlighting its potential for real-time rendering on low-capacity devices. Furthermore, OMG accelerates training speed. The substantial improvement over LocoGS can be attributed to two key factors: the reduced number of Gaussians and the absence of a large neural field. By efficiently exploiting coarse spatial information through a tiny MLP, OMG achieves high computational efficiency.
Table 3: Efficiency comparison of OMG variants evaluated on the Mip-NeRF 360 dataset. We present training time, the number of Gaussians, and the storage requirement with rendering quality.
@@
Mini-Splatting 19m 25s 531K 119.5 27.39 0.822 0.216
LocoGS-S
$\approx$1h
1.09M 7.9 27.04 0.806 0.232
@@
Figure 5: Evaluation without attribute compression.
Table 4: Ablation study of OMG using the Mip-NeRF 360 dataset.
@@
w/o SVQ 27.06 0.809 0.241 0.43M 19.8
@@ 4.3 Ablation study
Local distinctiveness (LD) scoring. OMG improves Gaussian pruning by incorporating LD scoring into the importance estimation. Without attribute compression, LD scoring enables high rendering quality with extremely reduced Gaussians, achieving similar performance compared to Mini-Splatting that use 20–30% more Gaussians, as shown in Figure˜5. When applying attribute compression (Table˜3), OMG with LD scoring still outperforms, leading to a significant performance improvement with a similar number of Gaussians. This effect becomes even more pronounced when the target Gaussian number is lower, demonstrating that LD scoring provides an effective approach for further reducing a sparse set of Gaussians.
OMG architecture. OMG leverages a highly compact neural field to capture coarse spatial information while reducing the number of learnable parameters per Gaussian. Section˜4.2 validates the contribution of this space feature. Although the total number of Gaussians slightly increases, performance significantly degrades. The absence of spatial information introduces instability in attribute learning, hindering effective importance scoring. This trend is consistently observed in both our small and medium models, highlighting the effectiveness of the space feature despite its minimal parameter overhead. Furthermore, when both the space feature and LD scoring are removed, the model experiences the most substantial performance drop. This indicates that the two contributions are orthogonal, independently contributing to model efficiency and performance.
Table 5: Ablation study of SVQ on the Mip-NeRF 360 dataset. We replace SVQ with either VQ or RVQ. ‘CB Init’ denotes the time (in seconds) required for codebook initialization.
Method OMG-XS OMG-M OMG-XL
[…]
OMG 4.8 4.06 27.06 7.6 5.31 27.21 7.9 6.82 27.34
[…]
Generalization ability. The OMG attribute representation is applicable to all 3DGS-based methods, as it flexibly represents attributes with spatial locality (via space features and SVQ) as well as those without (using SVQ alone). To validate this generalizability, we applied OMG to 3DGS-MCMC [mcmc], a method well-known for effective densification. As reported in Table˜6, OMG preserves the original performance and reduce storage requirements, demonstrating its broader applicability.
Table 6: Results of applying OMG representation to 3DGS-MCMC [mcmc] on the Mip-NeRF dataset.
Method #Gauss PSNR SSIM LPIPS Size #Gauss PSNR SSIM LPIPS Size
MCMC [mcmc] 500K 27.42 0.807 0.248 115 MB 1M 27.83 0.823 0.221 230 MB
MCMC+OMG 500K 27.21 0.797 0.256 5.1 MB 1M 27.63 0.813 0.227 10.0 MB
Table 7: Performance evaluation using Zip-NeRF [zipnerf] dataset. * indicates results reported in the SMERF [smerf] paper, where 3DGS hyperparameters were tuned for higher performance. We train and report results for 3DGS and Mini-Splatting using their default settings.
Method PSNR SSIM LPIPS #Gauss Size
Zip-NeRF* [zipnerf] 27.37 0.836 0.305 - 607 MB
SMERF* [smerf] 27.28 0.829 0.350 - 4.1 GB
3DGS* [3dgs] 25.50 0.810 0.369 - 212 MB
3DGS [3dgs] 25.16 0.813 0.358 934K 210 MB
Mini-Splatting [msv1] 24.57 0.802 0.370 337K 75.7 MB
MCMC [mcmc] 26.37 0.838 0.318 3M 675 MB
MCMC+OMG 26.69 0.839 0.310 3M 29.0 MB
5 Limitations and broader applicability
Baseline dependancy. OMG builds upon 3DGS and Mini-Splatting with the goal of preserving original rendering quality under compact representations. However, its applicability remains inherently bounded by the capabilities of these baselines. For instance, 3DGS and subsequent approaches exhibit clear limitations in capturing the complexity of larger-scale scenes such as the Zip-NeRF [zipnerf] dataset. As shown in Table˜7, both 3DGS and Mini-Splatting fail to densify a sufficient number of Gaussians, leading to degraded performance. Consequently, OMG inherits these limitations, restricting its applicability in such scenarios.
Broader applicability. On the other hand, as demonstrated in Table˜6, OMG can be applied in a generalizable manner across different baseline methods. Based on this strength, we tune 3DGS-MCMC for large-scale scenes and incorporate the OMG representation, achieving superior performance while drastically reducing storage requirements (Table˜7). More broadly, when integrated with ongoing research across diverse domains, OMG has the potential to provide a highly general and efficient representation.